OpenStudy (anonymous):

An anhydrous sample of calcium sulphate weighs 2.59 g. When mixed w/ water it hardens to form 3.54 g. of plaster of paris (CaSO4.xH20) a. Define x B. What is the formula for the hydrate? c. Name the hydrate

OpenStudy (anonymous):

\[ CaSO_4+xH_2O\longrightarrow CaSO_4\cdot xH_2O 1:x\qquad 1:x \] compare the number of moles

OpenStudy (anonymous):

follow? molar mass of CaSO4 \(= 40 + 32 + (16)(4) =136\) similarly, molar mass of H2O \(=2(1)+16=18\) molar mass of plaster of paris = \(136+18x\) in the problem, number of moles of CaSO4 \(\displaystyle = {2.59\over136}=0.019\) number of moles of plaster of paris = \(\displaystyle {3.54\over 136+18x}=0.019\) now, solve for "x"

OpenStudy (anonymous):

\[\Huge \color{Lime}{~\:\:\:\:\:\:\:\:\:\:\:\:\frak Welcome\:\:\frak To\:\:\frak Open\frak Study}\]

OpenStudy (anonymous):

Thanks!